Use the Dell Technologies, Inc. Form 10-K for the fiscal year ended February 1, 2019 to answer the computation questions. Use the stock price of $51.32, which is the closing stock price at June 14, 2019 for the stock price. Use 2.125% as the risk free rate. Use 5% as the market risk premium. Use 6% as the yield to maturity on the long-term debt. Use 1.08 as the beta. The financial statements may be found after the Report of Independent Registered Public Accounting Firm. Total assets on the balance sheet should be $111,820 million at February 1, 2019. Long-term debt amount should be $49,201 million at February 1, 2019. The following computations should be made for Dell Technologies, Inc.: Market capitalization Cost of equity Cost of debt Weighted average cost of capital

Unable to post the 10-K form because of the file size. I'm not sure what portion is required.
Answer #1

Interpretation of Question:

Information Provided :

Current market price = $51.32,

Risk free rate = 2.125%

Market risk premium = 5%

YTM (Long term debt) = 6 %

Beta =1.08

Long term debt = $49,201 million

Total Assets = $111,820 million

What to compute:

  1. Cost of equity
  2. Cost of debt
  3. WACC (Weightage Average Cost of Capital)

Perquisite Knowledge and Answer:

Meaning of cost of Capital:

Cost of capital is the rate of return expected by the existing capital providers.

There are two types of cost

  1. Cost of Specific sources of capital (E.g. Cost of Equity, cost of Debt, etc)
  2. Weighted Average Cost of Capital.

1.Cost of equity

For calculating cost of equity there are many Approaches:

  1. SML (Security Market Line) or CAPM Approach (Capital Assets Pricing Model)
  2. Bond Yield plus Risk Premium Approach
  3. Dividend Growth Model
  4. Earning Price Ratio Approach
  5. Others

So, we must see from given data that what is already given to choose appropriate method for calculating cost of Equity.

Here, as per given data suitable approach is SML.

Formula for Cost of equity Using SML Approach

Ke= Rf + β (Rm-Rf)

Where Ke = Cost of Equity

              β = Beta of investment

              Rf = RiskFree Rate

              Rm = Expected Market Return

              Rm-Rf = Market Risk Premium

So ,Ke = Rf + β (Rm-Rf)    

                   = 2.125%+1.08 (5%)

Ke = 7.525 %

2.Cost of Debt:

Here YTM is given so we will take it as Pre-Tax Kd = Pre-Tax Cost of Debt = 6%

From your K-10 Statements if Tax rate is provided than must calculate post tax cost of debt as follows:

Post-Tax Kd = (Pre-Tax Cost of Debt)(1-t),

Where t = Tax rate

For calculations of WACC only post tax cost of debt will be used.

3. WACC

   WACC means Weighted Average Cost of Capital

WACC = (Proportion of Equity) (cost of equity) + (Proportion of Debt) (Cost of Debt) + (Proportion Of any other source of finance given in capital structure) (Cost of respective source)

In this case you need to find all sources given in Form 10-K of dell Technologies, Inc.

Then, find proportion of each and multiply it with respective cost of that source and do summation of all weightage costs for finding WACC.

Here long-term debt value is provided i.e. $49,201 million, from statement find the value of equity and follow aforesaid step for finding WACC.

As here I don’t have whole statement, I can’t calculate WACC.